• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 04:56:05 on February 7, 1999

    Entry number 12677

    keyword=Beam Time Accounting
    Hall A

    2/6/1999 1600-2400
    Experiment #: E95-001
    Shift Leader: J.W. Martin
    Shift Workers: A. Deur, S. Strauch, S. Malov

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    1 25 1 1 1 1 45 1 7.16666666666667
    Beam Available
    but Not in Use
    0 0 0 0 0 0 15 0 0.25
    Beam Not Available
    or Unacceptable
    0 35 0 0 0 0 0 0 0.583333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 1 1 1 45 1 7.75
    Planned Config-
    uration Change
    0 0 0 0 0 0 15 0 0.25
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 17:15 accelerator system check begins, run 1683 ended prematurely 17:45 NMR measurement completed. P_target = 0.332 17:50 accelerator system check is done; did raster check; beam back. 18:00 premature end of run 1684 for T. Black to examine hadron d10 trigger; run 1685 is also short to test his modification; modification had no effect, so switching back to normal. 18:25 noticed beam spot position Y>.3; called MCC to get it fixed. 19:30 HAPPEX showed charge asym=340 ppm; called MCC to tweek pockels cell. They made a change. We run HAPPEX again to check it. 20:00 HAPPEX showed charge asym=-439 ppm; called MCC and told them to "turn it back half way". We will do another HAPPEX run. 20:40 HAPPEX shows charge asym=-110 ppm; we are happy. 22:24 Target spin direction changed to 62.5 degrees. Pedestal run done, too.

    Data Entry by: J.W. Martin